# Quantifying Self-Awareness of Knowledge in Large Language Models

### v1

@@ 3.2 Approximate question-side effect
$\mathcal{A}(\phi(s_{M}))\approx\mathcal{A}(\phi(s_{Q},s_{M}))-\mathcal{A}(\phi(s_{Q}))$ (2)
@@
To provide a clearer theoretical ground, we formalize AQE as a special case of Shapley analysis [fryer2021shapley], which evaluates the impact of individual factors on the outcome. Specifically, AQE corresponds to the concept of marginal contribution—a metric that quantifies the separate contriubtion of a single factor.
@@
Computing $A(\phi(s_{Q}))$ is achieved by introducing a distinct model $\theta^{\prime}$ (where $\theta^{\prime}\neq\theta$) which is optimized to only embed basic properties of the input question (e.g., domain or question type), $s^{\prime}_{Q}\approx s^{\prime}=\theta^{\prime}(x)$. A representative example of $\theta^{\prime}$ is sBERT [31]. sBERT is a very small model with only 22.7M parameters, but it is optimized to generate an embedding vector $s^{\prime}$ from input text $x$ (e.g., question). sBERT is known to capture high-level information from text as effectively as $\theta$ with a larger architecture (e.g., LLaMA-3-8B), as long as the target information is simple enough (e.g., domain classification). Therefore, while $s_{Q}^{\prime}$ and $s_{Q}$ reside in different representational spaces of two distinct models, they are assumed to capture similar high-level information ($s_{Q}^{\prime}\sim s_{Q}$). Conversely, due to its small size, we can assume $\theta^{\prime}$ holds a very small amount of knowledge, which makes the knowledge distribution of $\theta^{\prime}$ and $\theta$ independent ($s^{\prime}_{M}$ and $s_{M}$ are independent). This makes $s^{\prime}_{M}$ ignored when $\phi$ learning to predict $k$ (correctness of $\theta$) from $s^{\prime}$ (hidden state from $\theta^{\prime}$).
This results in following equation, where $\phi$ and $\phi^{\prime}$ share architecture but are trained independently, as $s_{Q}$ and $s_{Q}^{\prime}$ lie in different representational spaces.
@@
3.3 Measuring AQE across datasets
In this section, we measure AQE across hallucination prediction datasets. The model $\theta$ is LLaMA-3-8B-Instruct [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B, and we evaluate it mainly on short-form factoid datasets (e.g., ParaRel, Mintaka, Halueval, HotpotQA), and additionally on long-form factoid datasets (e.g., Explain). The details of each dataset are provided in the §D. We use two metrics: AUROC and accuracy. AQE for each metric is denoted as AQEauc and AQEacc, respectively. We also report the $p(k=True)$ and $p(k=False)$, the bias of binary label $k$. We present results for a larger (70B) model in §E, which shows similar trends.
As shown in Table 2, the AQEauc for most datasets is close to or exceeds 0.70, rather than centering around 0.5. This indicates that many datasets exhibit a strong effect of question-side shortcuts, and a model can easily achieve AUROC over 0.70 without any self-awareness (i.e., perception of model-side information), relying solely on question-aware skills such as domain classification. Though previous works [35, 42] have reported to achieve AUROC over 0.80 in hallucination prediction using these datasets, it could be a statistical artifact that is hard to generalize. In all datasets, AQEacc is consistently higher than the $p(k=True)$, indicating that AQE captures question-side effect beyond just the bias of the binary label.
3.4 AQE in refined dataset
In §3.1, we analyze that the type and the domain of a question can act as question-side shortcuts in predicting $k$. Fortunately, some datasets (HotpotQA, Mintaka, and ParaRel) include tags for this information, which allows us to control for it. We analyze AQE for each dataset before and after this control. In Table 9, the “+type” column refers to the dataset after excluding binary-type questions. “+domain” indicates a regrouping of the train and test data according to their categories, such that the domains do not overlap (i.e., out-of-domain setting). The refinement process is detailed in the §D.
@@ 3.4 AQE in refined dataset
The experimental results can be summarized in two main points. (1) Applying refinement leads to a significant reduction in AQE. This demonstrates that through post-processing, we can get a dataset with lower dependency on the question-side information, which is more suitable for evaluating self-awareness. (2) AQE still remains even after refinement. This suggests the presence of a question-side effect that we have not yet identified or controlled for.
4 SCAO: an approach to enhance the use of $s_{M}$
Causal LLM is an entity retriever when compressed The confidence score is calculated by the maximum inner product score between the last hidden state of $\theta$ and the token embedding vectors (i.e., decoder head). We analyze that this structure is analogous to the maximum inner product search (MIPS) used in dense retrieval [16] (Figure 3). And we focus on the calibration function of similarity scores in dense retrieval. Previous research on dense retriever systems such as Faiss [7] leverages this calibration through a range search, which finds all the document vectors that are within some distance threshold.
@@ 5 Experiment on hallucination prediction approaches
Approaches Previous approaches for hallucination prediction can be broadly categorized into three. (1) confidence-based: this utilizes the softmax probability of the answer token [9]. It is utilized in other forms, such as perplexity [32] or energy [23]. We adopt a simplified method that takes the mean of top-$n$ softmax probabilities of the first answer token and applying a threshold. $n$ and the threshold $t$ are learnable $\phi$. (2) hidden-state-based: This approach feeds the hidden-state vectors of a model $\theta$ into learnable $\phi$. $\phi$ can be a linear layer [19] or more complex architecture [1, 6]. We adopt a three-layer deep neural network. (3) aggregation: This approach concatenates the confidence scores and hidden state into a single vector, which is then passed to a learnable $\phi$ [35]. In Table 10, Conf, Probe, and Conf + Probe represent the three evaluated approaches, respectively. Detailed explanations are provided in §E.1. We conduct experiments using instruction-tuned LLaMA models of two different sizes (8B, 70B).
@@
(7) Even in a long-form question answering scenario (Explain), similar trends are observed, including the superiority of SCAO. This suggests that the “answer in one word” instruction remains effective even in scenarios where a one-word answer is inherently infeasible.
6 Conclusion and limitation
In this work, we provide a measurable definition of self-awareness in LLMs, distinguishing it from mere hallucination prediction by isolating model-side and question-side effects. We introduce AQE to quantify the question-side cues, enabling a clearer attribution of self-awareness. This definition and measurement offer a reference point for studying human-like mental traits in LLMs. Empirically, we show that much of the prior performance in hallucination prediction is driven by question-side shortcuts rather than self-awareness, and that our proposed method, SCAO, enhances self-awareness and yields more robust and generalizable predictions across diverse settings.
However, applying our framework beyond knowledge-recall tasks to more complex settings (e.g., reasoning, text reading, and long-form open-ended questions) remains an important direction for future work.
Appendix A Related works
Self-awareness in humans: Insights from cognitive neuropsychology
Extensive research in cognitive psychology and neuroscience has shown that human self-awareness—particularly in the context of knowing whether one knows something—relies on layered cognitive processes. According to studies such as [17, 13, 3], two key mechanisms underpin this self-assessment.
1) Unconscious level: When a query is received, the brain initiates implicit memory retrieval, evaluating whether candidate memories fit the contextual cues. This process activates regions such as the orbitofrontal and prefrontal cortices within 300–500 milliseconds [33, 13], distinct from areas responsible for linguistic output, such as the posterior temporal lobe (i.e., Wernicke’s area) [2]. Several factors—such as the amount, accessibility, and vividness of retrieved information—act as internal cues signaling potential knowledge [17].
The dual-process theory [15] offers a broader framing: rapid, intuitive processes dominate simple recall tasks, while complex, deliberative processes support tasks such as reasoning or problem-solving.
Self-awareness in LLMs: A perspective on hallucination detection
1) Before-generation: Approaches that attempt to detect potential hallucinations *before* the model generates an answer mirror the unconscious processes of humans. These methods, including our own, rely on internal indicators such as activation patterns or uncertainty proxies to determine knowledge sufficiency prior to verbalization [25]. Benchmarks like Mintaka [34] and ParaRel [8] primarily test immediate factual recall, aligning well with this layer of self-assessment.
2) After-generation: Other approaches evaluate the model’s response after generation, resembling conscious-level reasoning in humans. These include multi-pass generation, self-consistency checks, and the integration of external tools such as retrievers [5, 26, 6]. Benchmarks requiring structured reasoning, such as MMLU [12], TruthfulQA [22], and ELI5 [10], are more aligned with this stage, as they demand deliberative thought processes rather than pure retrieval.
Importantly, not all hallucinations can be attributed to failures in self-awareness. For instance, hallucinations in open-book QA tasks may arise from comprehension or inference errors rather than epistemic uncertainty. Benchmarks like SQuAD [30] and FEVER [36] exemplify this issue, being more about information grounding than self-assessment.
@@ Appendix B Other question-side shortcuts
Time-sensitive questions are inherently difficult for LLMs to answer accurately, as LLMs lack a robust understanding of time [14]. As a result, questions involving temporal information will be biased toward hallucinated responses. Datasets such as HaluEval, Mintaka, and TruthfulQA include such questions (e.g., "How old is Barack Obama?", "When did the most recent pandemic occur?").
@@ C.1 Efficiency of first token as a discriminator
Previous works on confidence-based hallucination detection research mostly utilize the confidence score of all tokens in the answer sentences, with normalization such as averaging [6]. Utilizing more information is ultimately more advantageous; however, it also has several drawbacks. We observe a pattern that as the entity name length increases, the average confidence tends to rise. For example, Figure 5 depicts the confidence pattern of the hallucinated question-answer pair “Question: Give me an explanation about Obama. Answer: Harry Potter and the Philosopher’s Stone”.
@@ D.1 Datasets and their refinement strategies
In this paragraph, we present details about the benchmark dataset for evaluating the hallucination prediction method: Mintaka [34], ParaRel [8], HaluEval [18], HotpotQA [39], and Explain. We also describe the refinement strategies applied to each dataset to reduce their AQE. “+ type” refers to refinements related to question types, while “+ domain” refers to refinements based on question domains, following Table 10
@@
ParaRel is originally a dataset designed for masked language modeling, containing factual knowledge expressed through diverse prompt templates and relational types. We utilize the rearranged version by [42]. This version consists of 25,133 prompt-answer pairs across 31 domains. It is further divided into two parts: the first 15 domains are classified as in-domain data, and the remaining 16 domains are classified as out-of-domain.
@@
HotpotQA is a question-answering dataset where each instance consists of a question, label (types including entity, boolean, numerical), and reference documents. We utilize only the question and answer to fit the closed-book scenario. An example of the question is “What government position was held by the woman who portrayed Corliss Archer in the film Kiss and Tell?”, paired with the label “Chief of Protocol”. We use the development dataset as a test set, following [42].
@@
We present a benchmark Explain to evaluate a model’s ability to provide a descriptive answer to an open-ended question. Explain is an extended and refined version of an open-ended long-form dataset in the well-known and verified work of FActScore [27]. In FActScore, a small dataset is devised to test the fact-checking pipeline for long-form QA. This dataset is created by appending prompts like “Tell me a bio of <entity>” to person names sourced from Wikipedia. However, its subjects are limited to only person names, and it includes only 500 entries.
@@ E.1 Detail on the three main approaches
(2) hidden-state-based: We employ a 3-layer deep neural network (DNN) structure with dimensions $d\rightarrow 100\rightarrow 30\rightarrow 1$, where $d$ is the hidden size. ReLU activation is applied between each layer. The objective function of DNN is binary cross entropy loss $L=-\frac{1}{N}\sum[y\cdot\log(\phi(s))+(1-y)\cdot\log(1-\phi(s))]$. DNN ($\phi$) is trained on the dataset while $\theta$ is frozen. The choice of which layer’s hidden state from $\theta$ to use is determined during the training phase, based on the one that achieves the highest validation accuracy. This approach extends the linear probing [19].
@@
(3) aggregation: We utilize the feature fusion of confidence and hidden state. That implies utilizing both top-30 confidence value and $h_{th}$ hidden state from $\theta$ as inputs to DNN. This approach concatenates the confidence scores and hidden state into a single vector, which is then passed to a learnable $\phi$ [35]. We use a module with dimensions of $(d+n)\rightarrow 100\rightarrow 30\rightarrow 1$, where $n$ is fixed to 30 (Figure 8).

### v2

@@ 4.2 Approximate question-side effect
$\mathcal{A}(\phi(s_{M}))\approx\mathcal{A}(\phi(s_{Q},s_{M}))-\mathcal{A}(\phi(s_{Q}))\vskip-20.0pt$ (2)
@@
We formalize AQE as a special case of Shapley analysis (Fryer et al., 2021), which evaluates the impact of individual factors on the outcome. Specifically, AQE corresponds to the concept of marginal contribution—a metric that quantifies the separate contribution of a single factor.
@@
Computing $A(\phi(s_{Q}))$ is achieved by introducing a distinct model $\theta^{\prime}$ (where $\theta^{\prime}\neq\theta$) which is optimized to only embed basic properties of the input question (e.g., domain or question type), $s^{\prime}_{Q}\approx s^{\prime}=\theta^{\prime}(x)$. A representative example of $\theta^{\prime}$ is sBERT (Reimers and Gurevych, 2019). sBERT is a very small model with only 22.7M parameters, but it is optimized to generate an embedding vector $s^{\prime}$ from input text $x$ (e.g., question). sBERT is known to capture high-level information from text as effectively as $\theta$ with a larger architecture (e.g., LLaMA-3-8B), as long as the target information is simple enough (e.g., domain classification). Therefore, while $s_{Q}^{\prime}$ and $s_{Q}$ reside in different representational spaces of two distinct models, they are assumed to capture similar high-level information ($s_{Q}^{\prime}\sim s_{Q}$). Conversely, due to its small size, we can assume $\theta^{\prime}$ holds a very small amount of knowledge, which makes the knowledge distribution of $\theta^{\prime}$ and $\theta$ independent ($s^{\prime}_{M}$ and $s_{M}$ are independent). This makes $s^{\prime}_{M}$ ignored when $\phi$ learning to predict $k$ (correctness of $\theta$) from $s^{\prime}$ (hidden state from $\theta^{\prime}$).
This results in Equation 4, where $\phi$ and $\phi^{\prime}$ share architecture but are trained independently, as $s_{Q}$ and $s_{Q}^{\prime}$ lie in different representational spaces.
@@
4.3 Measuring AQE across datasets
In this section, we measure AQE across hallucination prediction datasets. The model $\theta$ is LLaMA-3-8B-Instruct [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B, and we evaluate it mainly on short-form factoid datasets (e.g., ParaRel, Mintaka, Halueval, HotpotQA), and additionally on long-form factoid datasets (e.g., Explain). The details of each dataset are provided in the §D. We use two metrics: AUROC and accuracy. AQE for each metric is denoted as AQEauc and AQEacc, respectively. We also report the $p(k=True)$ and $p(k=False)$, the bias of binary label $k$. We show that a larger (70B) model shows similar trends (§E).
As shown in Table 1, the AQEauc for most datasets is close to or exceeds 0.70, rather than centering around 0.5. This indicates that many datasets exhibit a strong effect of question-side shortcuts, and a model can easily achieve AUROC over 0.70 without any self-awareness (i.e., perception of model-side information), relying solely on question-aware skills such as domain classification. Though previous works (Snyder et al., 2024; Zhang et al., 2024) have reported to achieve AUROC over 0.80 in hallucination prediction using these datasets, it could be a statistical artifact that is hard to generalize. In all datasets, AQEacc is consistently higher than the $p(k=True)$, indicating that AQE captures question-side effect beyond just the bias of the binary label.
4.4 AQE in refined dataset
In §4.1, we analyze that the type and the domain of a question can act as question-side shortcuts in predicting $k$. Fortunately, some datasets (HotpotQA, Mintaka, and ParaRel) include tags for this information, which allows us to control for it. We analyze AQE for each dataset before and after this control. In Table 9, the “+type” column refers to the dataset after excluding binary-type questions. “+domain” indicates a regrouping of the train and test data according to their categories, such that the domains do not overlap (i.e., out-of-domain setting). The refinement process is detailed in §D.
The experimental results can be summarized in two main points. (1) Applying refinement leads to a significant reduction in AQE. This demonstrates that through post-processing, we can get a dataset with lower dependency on question-side information, which is more suitable for evaluating self-awareness. (2) AQE still remains even after refinement. This suggests the presence of a question-side effect that we have not yet identified or controlled for.
4.5 SCAO
We propose a method called Semantic Compression by Answering in One word (SCAO), which leverages model-side information more effectively by instructing the model to “answer in one word,” thereby improving the alignment of confidence values. We provide further explanation in §C.
@@ 5 Experiment on hallucination prediction approaches
Approaches Previous approaches for hallucination prediction can be broadly categorized into three. (1) confidence-based: this utilizes the softmax probability of the answer token (Fadeeva et al., 2024). It is utilized in other forms, such as perplexity (Ren et al., 2023) or energy (Liu et al., 2021). We adopt a simplified method that takes the mean of top-$n$ softmax probabilities of the first answer token and applying a threshold. $n$ and the threshold $t$ are learnable $\phi$. (2) hidden-state-based: This approach feeds the hidden-state vectors of a model $\theta$ into learnable $\phi$. $\phi$ can be a linear layer (Li et al., 2024a) or more complex architecture (Azaria and Mitchell, 2023; Chen et al., 2024). We adopt a three-layer deep neural network. (3) aggregation: This approach concatenates the confidence scores and hidden state into a single vector, which is then passed to a learnable $\phi$ (Snyder et al., 2024). In Table 10, Conf, Probe, and Conf + Probe represent the three evaluated approaches, respectively. Detailed explanations are provided in §E.1. We conduct experiments using instruction-tuned LLaMA models of two different sizes (8B, 70B).
@@
(7) The $\mathcal{A}(\phi(s_{M}))$ is very low in OOD setting of Explain, which reveals that the hidden-state-based approach has limited generalization in long-form question answering settings. such as Explain. This contradicts the reports from previous works (Snyder et al., 2024; Chen et al., 2024).
6 Conclusion
In this work, we argued that hallucination prediction performance is often inflated by question-side shortcuts rather than reflecting genuine self-awareness. To disentangle these effects, we introduced Approximate Question-side Effect (AQE) and showed that many existing benchmarks exhibit strong question-aware signals, limiting the generalizability of prior results. Our analysis demonstrates that, once such effects are controlled, the contribution of true model-side awareness is substantially smaller than previously reported. Limitations are discussed in §A.
Appendix A Limitation
Scope Narrowed to System 1
As discussed in the §2, our study focuses on System 1 (fast, automatic processing) rather than System 2. The knowledge recall setting we consider largely involves rapidly retrieving stored information in response to a prompt; accordingly, our analyses and conclusions should be interpreted primarily as describing phenomena that arise in retrieval/recall-driven, short-horizon judgments and answer generation.
In contrast, tasks dominated by System 2 (slow, deliberative processing)—such as multi-step reasoning, planning, long-form generation, and explicit verification—may exhibit qualitatively different sources and signals of failure. Therefore, the indicators and observations proposed in this work should not be assumed to directly generalize to System 2–heavy settings.
That said, this scope choice does not diminish the significance of our contribution. Because System 1 and System 2 reflect fundamentally different computational regimes and error profiles, concepts like self-awareness and hallucination detection are likely to carry distinct meanings and objectives in each regime. Our work provides a foundation for more precise definition, measurement, and analysis of these concepts in System 1 settings, and clarifies what may need to be preserved versus redesigned when extending to System 2 scenarios.
Devising robust methods
How to design a more generalizable methodology remains an open question for future work. Although we propose SCAO as a methodology, it still exhibits a very high AQE (i.e., low genuine awareness) in long-form question answering. This indicates that hallucination prediction results for long-form QA should be interpreted with great caution when they are based on existing confidence- or hidden-state–based methods.
Moreover, this suggests that long-form answering involves more complex functions beyond simple knowledge recall, and therefore requires hallucination prediction approaches that go beyond solely leveraging a model’s internal states.
@@ Appendix B Other question-side shortcuts
Broken question The most commonly observed problem across all datasets is insufficient annotations of question and answer pairs. This occurs when questions and answers follow a one-to-many relationship, but the annotations fail to cover them. For example, the ParaRel consists of question-answer pairs such as: “Q: What field does Daniel Bernoulli work in? A: physics”. Although Daniel Bernoulli also worked in the fields of mathematics and medicine (one-to-many relation), only one label is provided, failing to cover all possible correct answers. If an LLM responds with a correct answer but different from the given label (e.g., mathematics), it would still be classified as hallucinated (i.e., $k$ is annotated $False$), which is incorrect. If $\phi$ is trained to predict $k$ using such $\{s,k\}$ pairs, it will likely become biased toward predicting $\hat{k}=False$ whenever the domain of question is “field of work”. This means $\phi$ learns a domain classification task, not self-awareness. While this may improve prediction performance within the dataset, it lacks generalizability. It is because its performance is likely to drop only if the quality of the dataset improves, or if questions are given from unseen domains. As a study of Zhang et al. (2024) reports hallucination performance on this dataset, this shortcut might have influenced the reported scores. This issue is also found in other datasets such as SimpleQuestions, which includes instances like “Q: What is a TV action show? A: Genji Tsushin Agedama”, that also fail to cover various possible answers.
Domain Classifying the domain of a question (e.g., science, society) alone can provide a rough guess of $k$. For example, suppose a model that is extensively trained on science data but is unfamiliar with society or history domains. Since the model is likely to make more hallucination when given questions in the history domain. In this way, $k$ can be biased toward $True$ or $False$ by domains, as shown in Figure 2. In such cases, the task of predicting $k$ with $\phi$ is more a domain classification, which is question-aware and not model-aware.
(a) HotpotQA
Type $p(k=True)$
Bridge 0.6828
Comparison 0.8477
(b) Mintaka
Type $p(k=True)$
Entity 0.5508
numerical 0.4011
date 0.5968
string 0.6315
boolean 0.7283
Question type Question type (e.g., short-answer, multiple-choice) also provides a strong hint for predicting $k$. The average probability of $k=True$ (denoted as $p(k=True)$) for binary-choice questions is at least 0.5 for random choice, significantly higher than for open-ended questions with 0 for random choice, as described in Table 4. In such a case, $\phi$ may learn a shortcut where it automatically predicts $k=True$ whenever it detects a binary-choice question. HotpotQA, HaluEval, and Mintaka contain such binary-choice questions.
@@
Time-sensitive questions are inherently difficult for LLMs to answer accurately, as LLMs lack a robust understanding of time (Jain et al., 2023). As a result, questions involving temporal information will be biased toward hallucinated responses. Datasets such as HaluEval, Mintaka, and TruthfulQA include such questions (e.g., "How old is Barack Obama?", "When did the most recent pandemic occur?").
@@ Appendix C Details on SCAO
Causal LLM is an entity retriever when compressed The confidence score is calculated by the maximum inner product score between the last hidden state of $\theta$ and the token embedding vectors (i.e., decoder head). We analyze that this structure is analogous to the maximum inner product search (MIPS) used in dense retrieval (Karpukhin et al., 2020) (Figure 3). And we focus on the calibration function of similarity scores in dense retrieval. Previous research on dense retriever systems such as Faiss (Douze et al., 2024) leverages this calibration through a range search, which finds all the document vectors that are within some distance threshold.
@@ C.1 Efficiency of first token as a discriminator
Previous works on confidence-based hallucination detection research mostly utilize the confidence score of all tokens in the answer sentences, with normalization such as averaging (Chen et al., 2024). Utilizing more information is ultimately more advantageous; however, it also has several drawbacks. We observe a pattern that as the entity name length increases, the average confidence tends to rise. For example, Figure 5 depicts the confidence pattern of the hallucinated question-answer pair “Question: Give me an explanation about Obama. Answer: Harry Potter and the Philosopher’s Stone”.
@@ D.1 Datasets and their refinement strategies
In this paragraph, we present details about the benchmark dataset for evaluating the hallucination prediction method: Mintaka (Sen et al., 2022), ParaRel (Elazar et al., 2021), HaluEval (Li et al., 2023), HotpotQA (Yang et al., 2018), and Explain. We also describe the refinement strategies applied to each dataset to reduce their AQE. “+ type” refers to refinements related to question types, while “+ domain” refers to refinements based on question domains, following Table 10
@@
ParaRel is originally a dataset designed for masked language modeling, containing factual knowledge expressed through diverse prompt templates and relational types. We utilize the rearranged version by (Zhang et al., 2024). This version consists of 25,133 prompt-answer pairs across 31 domains. It is further divided into two parts: the first 15 domains are classified as in-domain data, and the remaining 16 domains are classified as out-of-domain.
@@
HotpotQA is a question-answering dataset where each instance consists of a question, label (types including entity, boolean, numerical), and reference documents. We utilize only the question and answer to fit the closed-book scenario. An example of the question is “What government position was held by the woman who portrayed Corliss Archer in the film Kiss and Tell?”, paired with the label “Chief of Protocol”. We use the development dataset as a test set, following (Zhang et al., 2024).
@@
We present a benchmark Explain to evaluate a model’s ability to provide a descriptive answer to an open-ended question. Explain is an extended and refined version of an open-ended long-form dataset in the well-known and verified work of FActScore (Min et al., 2023). In FActScore, a small dataset is devised to test the fact-checking pipeline for long-form QA. This dataset is created by appending prompts like “Tell me a bio of <entity>” to person names sourced from Wikipedia. However, its subjects are limited to only person names, and it includes only 500 entries.
@@ E.1 Detail on the three main approaches
(2) hidden-state-based: We employ a 3-layer deep neural network (DNN) structure with dimensions $d\rightarrow 100\rightarrow 30\rightarrow 1$, where $d$ is the hidden size. ReLU activation is applied between each layer. The objective function of DNN is binary cross entropy loss $L=-\frac{1}{N}\sum[y\cdot\log(\phi(s))+(1-y)\cdot\log(1-\phi(s))]$. DNN ($\phi$) is trained on the dataset while $\theta$ is frozen. The choice of which layer’s hidden state from $\theta$ to use is determined during the training phase, based on the one that achieves the highest validation accuracy. This approach extends the linear probing (Li et al., 2024a).
@@
(3) aggregation: We utilize the feature fusion of confidence and hidden state. That implies utilizing both top-30 confidence value and $h_{th}$ hidden state from $\theta$ as inputs to DNN. This approach concatenates the confidence scores and hidden state into a single vector, which is then passed to a learnable $\phi$ (Snyder et al., 2024). We use a module with dimensions of $(d+n)\rightarrow 100\rightarrow 30\rightarrow 1$, where $n$ is fixed to 30 (Figure 8).